alice invested a portion of $32,000 at 9% interest and the balance at 11% interest. How much did she invest at each rate if her total income from both investments wa $3200.
Answer by checkley77(12844)

.11X+.09(32000-X)=3200
.11X+2880-.09X=3200
.02X=3200-2880
.02X=320
X=320/.02
X=16,000 INVESTED @ 11%.
32,000-16,000=16,000 INVESTED @ 9%
PROOF:
.11*16,000+.09*16,000=3200
1760+1440=3200
3200=3200